OpenStudy (anonymous):

A hyperbola is given by x2 − 4xy + y2 + 2 = 0. What is its equation in standard form if the axes are rotated so that the axes of the hyperbola lie on the rotated coordinate axes?

OpenStudy (anonymous):

\[(x' ^2)/(((3\sqrt(2))/2)) - (y' ^2)/(((3\sqrt(2))/2)) = 1 \]

OpenStudy (anonymous):

\[(x' ^ 2)/2 - (y' ^ 2)/((2/3)) = 1 \]

OpenStudy (anonymous):

\[(x' ^ 2)/((1/2)) - (y' ^ 2)/((1/2)) = 1 \]

OpenStudy (anonymous):

\[(x' ^ 2)/4 - (y' ^ 2)/4 = 1 \]

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

Holy ****.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

I will show you how it is set up

undeadknight26 (undeadknight26):

I am guessing you posted this on answers.com too eh? xD

OpenStudy (anonymous):

you already know

undeadknight26 (undeadknight26):

Hahah ;) Power of google.

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

I trying to show how it is set up but...

undeadknight26 (undeadknight26):

Waz wrong?

undeadknight26 (undeadknight26):

@nincompoop @skullpatrol ;-;

OpenStudy (anonymous):

\[\frac{ x^2 }{ (\frac{ 3\sqrt{2} }{ 2}}-\]

OpenStudy (anonymous):

ok thats the first half of A

OpenStudy (anonymous):

\[\frac{ y^2 }{ (\frac{ 3 \sqrt{2}}{ 2 }) }=\]

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

\[\frac{ x^2 }{ (\frac{ 3\sqrt{2} }{ 2})}- \frac{ y^2 }{ (\frac{ 3 \sqrt{2}}{ 2 }) }= 1\]

OpenStudy (anonymous):

And thats A

OpenStudy (anonymous):

but nobody knows what im doing so....

undeadknight26 (undeadknight26):

You are too advanced ;-; I am just a 10th grade simpleton ;-;

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

@satellite73